OpenStudy (anonymous):

Need help understanding this problem (1-i)^2, next(1-i)(1-i) = 1-i-i+i^2 =1-i-i+(-1), the ans. is -2i, and I don't see it?

OpenStudy (mathmate):

This is an application of the algebraic identity (a-b)^2=a^2-2ab+b^2 Now replace a=1, b=i, you get (1-i)^2=1^2-2(1)(i)+i^2 =1-2i+i^2 But i is a special number such that i^2=-1, so the net result is (1-i)^2 = 1-2i-1 = -2i after putting 1-1=0

OpenStudy (anonymous):

thankyou, I need to work on this one.

OpenStudy (mathmate):

You're welcome! :)